OpenStudy (anonymous):

suppose M varies jointly as S and A . if S is triples and A doubled, what happens to M ?

OpenStudy (anonymous):

M varies jointly as S and A means M = kSA

OpenStudy (anonymous):

\[ M_i = kSA \\ M_f = k\cdot 3S \cdot 2A M_f / M_i = \frac{ k\cdot 3S \cdot 2A}{kSA} \]You want to find M_f/M_i

OpenStudy (anonymous):

\[ M_i = kSA \\ M_f = k\cdot 3S \cdot 2A \\ M_f / M_i = \frac{ k\cdot 3S \cdot 2A}{kSA} \]

OpenStudy (anonymous):

3*2= 6 times and that too wio has gievn clear explanation and working